# Microscopic study of the low-energy enhancement in the gamma-decay strength of 50V

### v1

@@ II.1 Model space and effective interaction
For the calculations we have employed the $M$-scheme shell model code KSHELL [29]. For each of the total angular momenta $J=0,1,\ldots,8$ and for both parities, we have calculated 200 levels totalling 3600 levels, for which there are a total of 960000 $E1$ transitions and 958400 $M1$ transitions. In practice we had to exclude some of the transitions due to numerical reasons [^1] In some cases, two different levels end up having the exact same energy – to the floating point precision corresponding to an energy of 1 keV – leading to a gamma energy of 0. We have decided to exclude these relatively few transitions., resulting in a total of 959834 $E1$ transitions and 938347 $M1$ transitions. Having a large number of transitions is crucial for extracting the GSF since it describes average transition probabilities. We have used the SDPFSDG-MU effective interaction, which comprises the sdsd-pfpf-sdgsdg model space [30], see fig. 1 for an overview of the model space and orbitals. In this model space, 50V have 15 valence protons and 19 valence neutrons. This large valence space allows us to construct the full set of $0\hbar\omega$ and $1\hbar\omega$ basis states which are needed to describe the natural and unnatural parity states respectively (see section II.2 for more details). It should be stressed that taking the full $1\hbar\omega$ model space is highly desirable to obtain reliable $E1$ strengths.
@@ VI The GSF of 50V: features and physical mechanisms
From the left panel of fig. 8, it is immediately apparent that the calculated total dipole GSF displays a low-energy enhancement from approximately 3 MeV and down to 0 MeV. We can also clearly see that the LEE is due to the $M1$ part of the GSF, as the $E1$ GSF gives a negligible contribution in this energy region. While the $E1$ GSF starts at very low values, it steadily increases as $E_{\gamma}$ increases. At approximately 3 MeV there is a visible $E1$ contribution to the total strength, and by approximately 7 MeV the $E1$ GSF is at the same magnitude as the $M1$ GSF. It might seem like the $E1$ GSF decreases after 7 MeV; however, remember from section V that the GSF at the highest energies cannot be completely trusted because of low statistics and lack of level density.
@@
In the Oslo method, it is usually assumed that dipole transitions are dominating and the contributions from higher multipolarities are negligible. This assumption is supported by angular-distribution measurements in this mass region [7, 8]. As shown in fig. 8, the $E2$ contribution is negligible and can safely be ignored, supporting that it is reasonable to assume dominance of dipole transitions for this nucleus in this energy region. In the following, we will investigate in detail the nature of the LEE for the case of 50V.
@@ Appendix A Deriving the reduced transition probability
The creation-annihilation pair is however not a spherical tensor operator so we have to convert them into one before we can use the Wigner-Eckhart theorem. We know that if $\bm{T}_{L_{1}}$ and $\bm{T}_{L_{2}}$ are two spherical tensor operators of rank $L_{1}$ and $L_{2}$ respectively, then their tensor product – which is also a spherical tensor operator – is denoted $\bm{T}_{L}$ whose $2L+1$ components are defined as (see for example [41, sec. 2.3])

### v2

@@ II.1 Model space and effective interaction
For the calculations we have employed the $M$-scheme shell model code KSHELL [41]. For each of the total angular momenta $J=0,1,\ldots,8$ and for both parities, we have calculated 200 levels totalling 3600 levels, for which there are a total of 960000 $E1$ transitions and 958400 $M1$ transitions. In practice we had to exclude some of the transitions due to numerical reasons [^1] In some cases, two different levels end up having the exact same energy – to the floating point precision corresponding to an energy of 1 keV – leading to a gamma energy of 0. We have decided to exclude these relatively few transitions., resulting in a total of 959834 $E1$ transitions and 938347 $M1$ transitions. Having a large number of transitions is crucial for extracting the GSF since it describes average transition probabilities.
We have taken the full sdsd-pfpf-sdgsdg shell as the valence shell, see fig. 1 for an overview of the model space and orbitals. In this model space, 50V have 15 valence protons and 19 valence neutrons. This large valence space allows us to construct the full set of $0\hbar\omega$ and $1\hbar\omega$ basis states which are needed to describe the natural and unnatural parity states respectively (see section II.2 for more details). It should be stressed that taking the full $1\hbar\omega$ model space is highly desirable to obtain reliable $E1$ strengths because the $E1$ operator produces $1\hbar\omega$ states when applied to $0\hbar\omega$ states. It is also noted that the level densities of unnatural-parity states are close to those of natural-parity states for $E_{x}\gtrsim 3$ MeV as shown later, thus playing an essential role also in the low-energy $M1$ transitions.
We have used the SDPFSDG-MU effective interaction [52]. This interaction is based on the USD interaction [6] for the $sd$-shell part, the GXPF1B interaction [15] for the $pf$-shell part, and a variant of the $V_{\textrm{MU}}$ interaction [33] that is employed in the SDPF-MU interaction [48] for the remaining part. In addition, a few two-body matrix elements are modified in the same way as the SDPF-MU interaction, see Ref. [48] for more details. Since the Fermi surface of 50V is located in the $pf$-shell, the $pf$-shell interaction is particularly important for describing this nucleus. The GXPF1B interaction was constructed in [15] to achieve a good description of neutron-rich Ca isotopes with minimal modifications to the semi-empirical GXPF1 interaction [13].
@@ VI The GSF of 50V: features and physical mechanisms
From the left panel of fig. 8, it is immediately apparent that the calculated total dipole GSF displays a low-energy enhancement from approximately 3 MeV and down to 0 MeV. We can also clearly see that the LEE is due to the $M1$ part of the GSF, as the $E1$ GSF gives a negligible contribution in this energy region. While the $E1$ GSF starts at very low values, it steadily increases as $E_{\gamma}$ increases. At approximately 3 MeV there is a visible $E1$ contribution to the total strength, and by approximately 7 MeV the $E1$ GSF is at the same magnitude as the $M1$ GSF. It might seem like the $E1$ GSF decreases after 7 MeV; however, remember from section V that the GSF at the highest energies cannot be completely trusted because of low statistics and low level density.
@@
In the Oslo method, it is usually assumed that dipole transitions are dominating and the contributions from higher multipolarities are negligible. This assumption is supported by angular-distribution measurements in this mass region [24, 26]. As shown in fig. 8, the $E2$ contribution is negligible and can safely be ignored, supporting that it is reasonable to assume dominance of dipole transitions for this nucleus in this energy region.
To test the impact of the chosen interaction and model space, we have also performed calculations with the KB3G interaction [34]. The same quenching of 0.9 as well as $g_{l}=1.1,-0.1$ was used. In fig. 9 we compare the $M1$ strength of KB3G (green) and SDPFSDG-MU (red). In the LEE region they are generally very similar, however the KB3G $M1$ strength is slightly higher. The discrepancy can be explained by the fact that KB3G uses the $pf$ shell as model space and thus only captures positive-to-positive parity transitions but not negative-to-negative parity transitions. On the other hand, the SDPFSDG-MU interaction extends over more than one major shell and captures both types of transitions, giving us a more complete picture of the $M1$ strength.
To illustrate the implications, in fig. 9, we have plotted separately the positive-to-positive and negative-to-negative parity $M1$ strengths of SDPFSDG-MU. The positive-to-positive LEEs of the two interactions are very similar in both shape and amplitude, which shows that the calculations are quite robust with respect to the chosen interaction. Interestingly, the negative-to-negative transitions contribute to a reduction in the LEE, an effect which is not captured by a one-major-shell interaction. In the following, we will investigate in detail the nature of the LEE for the case of 50V.
Figure 9: $M1$ strength functions calculated with the interactions SDPFSDG-MU and KB3G in red and green, respectively. In orange and dark red we have $M1$ strength functions with SDPFSDG-MU where only positive-to-positive and negative-to-negative parity transitions have been used, respectively.
@@ Acknowledgements.
The data that support the findings in this article are openly available at [8].
@@ Appendix A Deriving the reduced transition probability
The creation-annihilation pair is however not a spherical tensor operator so we have to convert them into one before we can use the Wigner-Eckhart theorem. We know that if $\bm{T}_{L_{1}}$ and $\bm{T}_{L_{2}}$ are two spherical tensor operators of rank $L_{1}$ and $L_{2}$ respectively, then their tensor product – which is also a spherical tensor operator – is denoted $\bm{T}_{L}$ whose $2L+1$ components are defined as (see for example [46, sec. 2.3])